# Second FRCSyn-onGoing: Winning Solutions and Post-Challenge Analysis to Improve Face Recognition with Synthetic Data

### v1

@@ Abstract
Figure 1: Examples of synthetic identities and variations for different demographic groups using GANDiffFace [21].
Synthetic data has gained popularity as a good solution to mitigate some of these drawbacks [22, 23], allowing the generation of i) a large number of facial images from different non-existent identities, and ii) variability in terms of demographic attributes and scenario conditions. Several Generative AI approaches have been presented in the last couple of years for the synthesis of face images, considering state-of-the-art deep learning methods. Two of the most popular methods are Generative Adversarial Networks (GANs) [24] and Diffusion Models [25]. Popular generative methods can be SynFace [26] and SFace [27], based on GANs, DCFace [28], IDiff-Face [29] and ID3PM [30], based on Diffusion models, GANDiff-Face [21], based on the combination of GANs and Diffusion models, or alternative methods such as DigiFace-1M [31] and ITI-GEN [32], based on computer graphics techniques and a vision-language pre-trained CLIP [33] model, respectively. Some examples of synthetic face images generated using GANDiffFace are shown in Figure 1.
Beyond the generation of synthetic faces, another critical aspect lies in understanding the potential applications and benefits of synthetic data in enhancing FR technology. Recent studies have highlighted a performance gap between FR systems trained only with synthetic data and those trained on real data [28, 26]. Nevertheless, the results achieved in the 1 edition of the Face Recognition Challenge in the Era of Synthetic Data (FRCSyn) [34, 35], emphasize the relevance of synthetic data, either alone or merged with real data, in mitigating challenges in FR, such as demographic bias [34, 35]. Notably, in the 1 FRCSyn-onGoing, only synthetic data from DCFace [28] and GANDiffFace [21] methods were allowed for training FR systems. Additionally, together with novel generative methods, improving FR technology involves refining the design and training processes to address domain gaps between real and synthetic data in certain scenarios. For instance, observations from the 1 FRCSyn-onGoing revealed that most teams considered similar deep learning architectures (e.g., ResNet-100 [36]) and loss functions (e.g., AdaFace [7]), commonly used in FR systems trained with real data.
@@ II Second FRCSyn-onGoing: Databases
TABLE I: Description of some possible generative methods and synthetic databases that can be used by participants in the 2 FRCSyn-onGoing.
Id = Identities, Img = Images
@@
TABLE II: Description of the real databases considered for evaluation in the 2 FRCSyn-onGoing.
Id = Identities, Img = Images
@@
TABLE III: Tasks and sub-tasks for the 2 FRCSyn-onGoing and their respective metrics and databases.
SD = Standard Deviation. ID = Identity
Task 1: synthetic data for demographic bias mitigation
Sub-Task 1.2: [unconstrained] training exclusively with synthetic data
Train: no restrictions in terms of the number of face images.
Eval: BUPT-BalancedFace.
Sub-Task 1.3: [constrained] training with real and synthetic data
Train: CASIA-WebFace, and maximum face synthetic images.
Eval: BUPT-BalancedFace.
Task 2: synthetic data for overall performance improvement
Baseline: training with only CASIA-WebFace.
Metrics: accuracy (for each evaluation database).
Ranking: average accuracy, see Section III-C for more details.
Sub-Task 2.1: [constrained] training with only synthetic data
Train: maximum face images.
Eval: BUPT-BalancedFace, AgeDB [44], CFP-FP [45], and ROF [46].
Sub-Task 2.2: [unconstrained] training with only synthetic data
Train: no restrictions in terms of the number of face images.
Eval: BUPT-BalancedFace, AgeDB, CFP-FP, and ROF.
Sub-Task 2.3: [constrained] training with real and synthetic data
Train: CASIA-WebFace, and maximum face synthetic images.
Eval: BUPT-BalancedFace, AgeDB, CFP-FP, and ROF.
@@ III Second FRCSyn-onGoing: Setup
Due to the success of the 1 FRCSyn-onGoing [34, 35], we also decided to run the 2 edition in Codalab111111https://codalab.lisn.upsaclay.fr/competitions/16970, an open-source framework designed for conducting scientific competitions and benchmarks. On this platform, participants can find the competition’s requirements and limitations and can submit their scores to automatically obtain i) the evaluation metrics of their system, and ii) the position on the challenge leaderboard.
@@
Similar to the 1 FRCSyn-onGoing [35, 34], in this 2 edition we also explore the application of synthetic data for training FR systems, with a specific focus on addressing two critical aspects in current FR technology: i) mitigating demographic bias, and ii) enhancing overall performance under challenging conditions that include variations in age and pose, the presence of occlusions, and diverse demographic groups. To investigate these two areas, we consider two different tasks, each comprising three sub-tasks. Each sub-task considers different types (real/synthetic) and amounts of data for training the FR systems. Consequently, the 2 edition comprises 6 different sub-tasks. We summarize in Table III the key aspects of the experimental protocol, metrics, and restrictions for each sub-task.
@@ IV Second FRCSyn-onGoing: Systems Description
In this 2 FRCSyn-onGoing, we encourage participants to propose novel Generative AI methods for the creation of synthetic data. Besides, we also give participants the freedom to choose the FR architecture and training methods. Table LABEL:tab:data_meth summarizes for each team the key information in terms of the proposed synthetic data and FR system. Teams are arranged by their average ranking in the 6 sub-tasks from the 2 FRCSyn-onGoing. In general, we can see that most teams have decided to use synthetic data from DCFace [28] and IDiff-Face [29] databases, improving also the original data through cleaning and selection approaches, among other more sophisticated techniques. Also, regarding the FR technologies, most of them are based on ResNet [36] and IResNet [48] architectures, with AdaFace [7] and ArcFace [6] as the main used losses. However, some teams proposed their own methods to generate synthetic facial images, as well as to train their FR models. Next, we describe the specific details of the top-11 proposed systems in the 2 FRCSyn-onGoing.
@@
This team comprises members from Fudan University and Tencent Youtu Lab, China. They used a Latent Diffusion Model (LDM) [49] based on IDiff-Face [29] to synthesize faces. The LDM was conditioned using identity embeddings as contexts, extracted from faces by a pretrained ElasticFace [50] model with an IResNet-101 [48] backbone. They trained the LDM with the CASIA-WebFace [42] database. As the LDM takes the ID embeddings as context, they considered an unconditional Denoising Diffusion Probabilistic Model (DDPM) trained on the FFHQ database [51] as a context generator. Specifically, they used the DDPM to generate faces with arbitrary identities, known as context faces. They exploited the same ElasticFace model to extract the embeddings from the context faces. To encourage the quality and the distinctiveness of identity from later LDM-generated faces, they filtered the embeddings by setting a minimum cosine similarity threshold of between arbitrary pairs of embeddings. This yields embeddings with discriminative identities. Furthermore, they accelerated the sampling process of the LDM by Denoising Diffusion Implicit Models (DDIM) [52]. For the training of the FR model, they generated images for each context. They adopted the ID oversampling strategy from DCFace [28] and performed it five times for each ID to enhance consistency. As a result, contexts were used for Sub-Tasks 1.1 and 2.1, while for Sub-Tasks 1.2 and 2.2. For Sub-Tasks 1.3 and 2.3, they expanded Sub-Tasks 1.1 and 2.1 with the CASIA-WebFace database. Both the baseline and proposed FR models used IResNet-101 architectures. They applied the ArcFace [6] loss with a batch size of and an initial learning rate of for epochs. The learning rate was divided by at epochs , , and . They also used random cropping augmentation during training. Their proposed architecture is described in Figure 2.
Figure 2: Framework proposed by the ADMIS team.Figure 3: Framework proposed by the OPDAI team.Figure 4: Framework proposed by the ID R&D team.
@@
This team comprises members from Interactive Entertainment Group of Netease Inc, China. They initially used the DCFace [28] database, generating then more face images for each ID with large pose variations and occlusions using Photomaker [53]. They randomly replaced these images in the original DCFace data to ensure that the total number of samples meets the requirement of . During the Photomaker inference, they adopted a batch size of and used random prompts including age, pose, and image quality to ensure the diversity of the generated samples. For Sub-Tasks 1.2 and 2.2, they combined this data with the version of DCFace, while for Sub-Tasks 1.3 and 2.3, it was merged with CASIA-WebFace [42]. For Sub-Tasks 1.2, 1.3, 2.2, and 2.3 they did not merge nor denoise samples from different databases, following the Partial FC approach [64]. Regarding the FR model, they obtained the loss of different databases in independent AdaFace [7] heads, calculating the final loss as the average of the multiple heads. Both baseline and proposed models are based on IResNet-100 [48] architectures, with horizontal flipping. Their proposed architecture is described in Figure 3.
@@
This team comprises members from ID R&D Inc, USA. To generate the synthetic data, they used two models trained on WebFace42M [65]. The first model was based on Hourglass Diffusion Transformers (HDT) [54]. It was trained focusing on conditional flow matching [66] and following the classifier-free guidance approach [67]. Identity embeddings were used directly, whereas style embeddings were processed through a Vector Quantised-Variational AutoEncoder (VQVAE) [55]. Specifically, for head position, embeddings were allocated for VQVAE processing, while age and facial expression attributes were represented with embeddings each. During inference, combinations of these embeddings were randomly selected. The second model used to generate synthetic data was based on StyleNAT [56], enhanced with a FR model. This network was trained using auxiliary sources of supervision: a pre-trained FR network with Prototype Memory [68] (for identity supervision) and a pre-trained face attribute classification network (for style supervision). To create their synthetic data, they used classifier weights of the trained Prototype Memory to get identity embeddings, of which were randomly selected and were uniformly sampled from the clusters obtained using k-means, to get demographic diversity. For each identity, they generated face images using each of the two generative models. In the first stage, identity embeddings were used by the HDT model to get images for each identity. Then of these images were included in the training dataset. Identity and style embeddings were taken from the remaining images and used as a condition to generate different images with the StyleNAT model. These images were also put into the training dataset. Regarding the FR model, they trained an IResNet-200 [48] with UniFace [58] loss for epochs. One network was trained with color, geometric augmentations, and FaceMix-B [57], and the other one using only random horizontal flipping. These two networks were combined in an “ensemble”, where the first one received the original image, and the second one a mirrored copy. They used the same model for Sub-Tasks 1.1, 1.2, 2.1 and 2.2. For Sub-Tasks 1.3 and 2.3, they combined the synthetic data with the CASIA-WebFace [42] data, training two models, one on the mixed data, and the other on the CASIA-WebFace. Their proposed architecture is described in Figure 4.
Figure 5: Framework proposed by the K-IBS-DS team.
@@
This team comprises members from Korea Advanced Institute of Science & Technology and Institute for Basic Science, South Korea. They used DCFace [28] with and face images (depending on the sub-task). Regarding the FR model, they used several IResNet [48] models of , , and layers with Squeeze-and-Excitation (SE) blocks [60]. Inspired by SlackedFace [59], they made two modifications to enhance the AdaFace [7] FR classifier: i) they used renormalized uniform initialization as a more reliable weight initialization for uniformity across identity prototypes in the unit sphere and ii) replaced the L2-norm with the powered-norm (p-norm), or face recognizability index from [59], which integrates the L2-norm with the learned embedding proximity. The training stage was in line with [28] and [31], including optimizer, learning rate, etc. For Sub-Tasks 1.3 and 2.3, the first subjects of the CASIA-WebFace [42] were assigned for training, and the remaining ones for performance validation using random pairs with challenging conditions (identified based on the poorest L2-norm values [21]). The final score was obtained by aggregating the comparison scores of the different IResNet models, along with the horizontally flipped instances through score fusion. All training and test sets were realigned using RetinaFace [69], followed by a similarity transformation. For all sub-tasks, aggressive data augmentations were applied, including random horizontal flipping, photometric operations, cropping, resizing, and the addition of sunglasses and masks. Their proposed architecture is described in Figure 5.
Figure 6: Framework proposed by the Idiap-SynthDistill team.
@@
This team comprises members from China Telecom AI, China. By analyzing popular synthetic data, they found that intra-class and inter-class noise was widely present. Data cleaning can effectively remove the bad examples of synthetic data and retain important images from a large amount of synthetic data. In order to select the optimal synthetic data, they first trained an IResNet-100 [48] model with SE blocks [60] using CASIA-WebFace [42] to extract features of synthetic images from DCFace [28], GANDiffFace [21], and DigiFace-1M [31]. Subsequently, they used DBSCAN clustering to segregate intra-class noise and removed IDs with a class center feature cosine similarity greater than . Finally, they used the cleaned synthetic data merged with CASIA-WebFace to finetune the IResNet-100 for a second data refinement. From the final refined synthetic dataset, they sampled face images while retaining as many IDs as possible to build their synthetic training set. Regarding the FR model, in particular Sub-Task 2.3 in which they achieved their highest position among all sub-tasks, they trained IResNet-100 with AdaFace loss [7] (A1) and CosFace loss [61] (A2) with mask and occlusion augmentation on CASIA-WebFace and the refined synthetic data. They used an ensemble of A1, A2, and a model trained with only synthetic data. Furthermore, data augmentation was considered to enhance all features.
@@
This team comprises members from Idiap Research Institute, EPFL, and Université de Lausanne, Switzerland. The proposed method was based on SynthDistill [62], which is an end-to-end approach, generating synthetic images and training the FR model in the same training loop. Instead of using the pre-trained model in a separate step, they directly used it in the training loop for supervision, while a new student FR model was trained fully using synthetic data generated from a StyleGAN model [51]. For generating synthetic images, they trained StyleGAN2 [70] with the CASIA-WebFace database [42] and then dynamically generated 20 synthetic images during training based on the training loss. For the dynamic image generation, they used the training loss from every iteration as feedback to find the most difficult synthetic image in each batch and then re-sampled a new batch of synthetic images in the intermediate latent space of StyleGAN near the latent vector of the most difficult sample. If the loss value was high, they re-sampled with a relatively small standard deviation around the difficult sample and generated similar images, but if the loss value was small they re-sampled with a higher standard deviation, generating images with more variations. Throughout the process, the generated images were resized to before being fed to the FR models. They used a pre-trained FR model with IResNet-101 [48] architecture trained with CosFace [61] loss on a subset of the WebFace260M database [65] and trained a new model as a student network with the same architecture using synthetic data with their dynamic synthetic image generation approach. They used the Adam optimizer with an initial learning rate of and trained their student model with the same loss function as in [62]. For thresholding, a subset of DCFace [28] was used to determine the optimal threshold for maximizing verification accuracy, using a 10-fold cross-validation approach based on a random selection of identities and comparison pairs. Their proposed architecture is described in Figure 6.
Figure 7: Architecture proposed by the INESC-IGD team.
@@
This team comprises members from Fraunhofer IGD, Germany, University of Cagliari, Italy, and University of Ljubljana, Slovenia. They used the DCFace [28] synthetic database as it led to remarkable performance gains under well-known evaluation benchmarks for face verification, while combined with real data [74]. Also, they considered synthetic data generated with IDiff-Face [29] and ExFaceGAN [63] in Sub-Tasks 1.1, 1.2, 2.1, and 2.2. The ExFaceGAN data was generated using an identity disentanglement approach on pretrained GAN-Control [75]. Regarding the FR model, they trained a ResNet-100 [36] network using CosFace loss [61] with a margin penalty of 0.35 and a scale term of 64. The similarity mean difference between real-only and synthetic-only samples was scaled and added to the loss value. They trained the FR model for epochs with a batch size of and an initial learning rate of , which was divided by after , , , and epochs. During the training phase, the synthetic samples were augmented using RandAugment with operations and a magnitude of 16, following [76, 74]. For Sub-Tasks 1.3 and 2.3, the selected synthetic dataset was combined with CASIA-Webface [42], obtaining a total of images from identities. Their proposed architecture is described in Figure 8.
Figure 8: Framework proposed by the UNICA-IGD-LSI team.
@@
Figure 9: Framework proposed by the SRCN_AIVL team.

### v2

@@ Abstract
Index Terms:
FRCSyn, Face Recognition, Synthetic Data, Generative AI, Demographic Bias, Benchmark, PrivacyI Introduction
@@
Synthetic data has gained popularity as a good solution to mitigate some of these drawbacks [24, 25], allowing the generation of i) a large number of facial images from different non-existent identities, and ii) variability in terms of demographic attributes and scenario conditions. In this context, we refer to demographics as whole societies or smaller groups defined by criteria such as ethnicity, sex, or age [11, 26]. Several Generative AI approaches have been presented in the last couple of years for the synthesis of face images, considering state-of-the-art deep learning methods. The most popular approaches to generate synthetic facial images are Generative Adversarial Networks (GANs) [27], Diffusion Models [28] or the combination of both [21]. However, there are other less common approaches that do not rely on these models and still achieve a high level of realism in image generation [29, 30]. Some examples of synthetic face images generated using some of these methods are shown in Figure 1.
Beyond the generation of synthetic faces, another critical aspect lies in understanding the potential applications and benefits of synthetic data in enhancing FR technology. Recent studies have highlighted a performance gap between FR systems trained only with synthetic data and those trained on real data [22, 31]. Nevertheless, the results achieved in the 1 edition of the Face Recognition Challenge in the Era of Synthetic Data (FRCSyn) [32, 33], emphasize the relevance of synthetic data, either alone or merged with real data, in mitigating challenges in FR, such as demographic bias [32, 33]. Notably, in the 1 FRCSyn-onGoing, only synthetic data from DCFace [22] and GANDiffFace [21] methods were allowed for training FR systems. Additionally, together with novel generative methods, improving FR technology involves refining the design and training processes to address domain gaps between real and synthetic data in certain scenarios. For instance, observations from the 1 FRCSyn-onGoing revealed that most teams considered similar deep learning architectures (e.g., ResNet-100 [34]) and loss functions (e.g., AdaFace [7]), commonly used in FR systems trained with real data. Moreover, the use of synthetic facial data is not limited only to FR. With the emerging popularity of foundational models [35, 36], synthetic facial data can also be leveraged to provide these large models with a general understanding of what a human face looks like, serving as pretraining for many other tasks [37, 38, 39]. Additionally, the fact that these data can be labeled as they are generated allows the creation of novel datasets for various tasks such as attribute detection, facial expression recognition, and more. Although in the field of synthetic faces there are still not many public datasets in these lines of research, several studies have already highlighted the benefits of synthetic data for biometric tasks such as facial expression recognition [40], signature verification [41], action recognition [42], or pose estimation [43].
@@
Database | Framework | # Id | # Img/Id | # Img | ||
DCFace [22] | Diffusion Model | 20K | 50 | 1,200K | ||
40K | 5 | |||||
GANDiffFace [21] | GAN | 10K | 50 | 500K | ||
Diffusion Model | ||||||
Diffusion Model | 10K | 50 | 500K | ||
Diffusion Model | 10K | 50 | 500K | ||
DigiFace-1M [29] | 3D-model | 10K | 72 | 1.2M | ||
100K | 5 | |||||
ID3PM [45] | Diffusion Model | - | - | - | ||
SFace [46] | GAN | 10K | 60 | 0.6M | ||
SYNFace [31] | GAN | 10K | 100 | 1M | ||
ITI-GEN [30] | CLIP | - | - | - |
@@ II-A Synthetic Databases
SFace888https://github.com/fdbtrs/SFace-Privacy-friendly-and-Accurate-Face-Recognition-using-Synthetic-Data [46]: This framework uses a conditional GAN to synthetically generate face images with an adaptive discriminator augmentation to increase the diversity of the training database.
@@
Id = Identities, Img = Images
TO = Trade-Off. GAP = Gap to Real. AVG = Average Verification Accuracy. SD = Standard Deviation. FLOPs = Floating Point Operations Per Second. SYN = Accuracy Proposed Model. REAL = Accuracy Baseline Model. ID = Identity.
Task 1: synthetic data for demographic bias mitigation | |
Baseline: training with only CASIA-WebFace [51]. | |
Ranking: Trade-Off, see Section III-C for more details. | |
Metrics: | |
Sub-Task 1.1: [constrained] training exclusively with synthetic data | |
Train: maximum face images (e.g., IDs and images per ID). | |
Eval: BUPT-BalancedFace [52]. | |
Sub-Task 1.2: [unconstrained] training exclusively with synthetic data | |
Train: no restrictions in terms of the number of face images. | |
Eval: BUPT-BalancedFace. | |
Sub-Task 1.3: [constrained] training with real and synthetic data | |
Train: CASIA-WebFace, and maximum face synthetic images. | |
Eval: BUPT-BalancedFace. | |
Task 2: synthetic data for overall performance improvement | |
Baseline: training with only CASIA-WebFace. | |
Ranking: average accuracy, see Section III-C for more details. | |
Metrics: | |
Sub-Task 2.1: [constrained] training with only synthetic data | |
Train: maximum face images. | |
Eval: BUPT-BalancedFace, AgeDB [53], CFP-FP [54], and ROF [55]. | |
Sub-Task 2.2: [unconstrained] training with only synthetic data | |
Train: no restrictions in terms of the number of face images. | |
Eval: BUPT-BalancedFace, AgeDB, CFP-FP, and ROF. | |
Sub-Task 2.3: [constrained] training with real and synthetic data | |
Train: CASIA-WebFace, and maximum face synthetic images. | |
Eval: BUPT-BalancedFace, AgeDB, CFP-FP, and ROF. | |
Restrictions: | FLOPS50 GFLOPS |
Only the specified databases can be used for training. | |
Generative models cannot be used to generate supplementary data. |
@@ III Second FRCSyn-onGoing: Setup
Due to the success of the 1 FRCSyn-onGoing [32, 33], we also decided to run the 2 edition in Codalab111111https://codalab.lisn.upsaclay.fr/competitions/16970, an open-source framework designed for conducting scientific competitions and benchmarks. On this platform, participants can find the competition’s requirements and limitations and can submit their scores to automatically obtain i) the evaluation metrics of their system, and ii) the position on the challenge leaderboard. Table III provides an overview of the key aspects of the experimental protocol, metrics and restrictions for each sub-task. More detailed explanations can be found in their respective subsections.
@@ III-A Tasks
Similar to the 1 FRCSyn-onGoing [33, 32], in this 2 edition we also explore the application of synthetic data for training FR systems, with a specific focus on addressing two critical aspects in current FR technology: i) mitigating demographic bias, and ii) enhancing overall performance under challenging conditions that include variations in age and pose, the presence of occlusions, and diverse demographic groups. To investigate these two areas, we consider two different tasks, each comprising three sub-tasks. Each sub-task considers different types (real/synthetic) and amounts of data for training the FR systems. Consequently, the 2 edition comprises 6 different sub-tasks.
@@ IV Second FRCSyn-onGoing: Systems Description
In this 2 FRCSyn-onGoing, we encourage participants to propose novel Generative AI methods for the creation of synthetic data. Besides, we also give participants the freedom to choose the FR architecture and training methods. Table IV summarizes for each team the key information in terms of the proposed synthetic data and FR system. This table serves as a quick reference, while more detailed explanations of each team’s approach and methodology can be found in the corresponding subsections. Teams are arranged by their average ranking in the 6 sub-tasks from the 2 FRCSyn-onGoing. In general, we can see that most teams have decided to use synthetic data from DCFace [22] and IDiff-Face [23] databases, improving also the original data through cleaning and selection approaches, among other more sophisticated techniques. Also, regarding the FR technologies, most of them are based on ResNet [34] and IResNet [57] architectures, with AdaFace [7] and ArcFace [6] as the main used losses. However, some teams proposed their own methods to generate synthetic facial images, as well as to train their FR models. Next, we describe the specific details of the top-11 proposed systems in the 2 FRCSyn-onGoing.
Team | Country |
Fig. |
Synthetic Data - Team Improvements | FR Model - Training method | ||||||||
ADMIS | China | All | 2 | Novel |
OPDAI | China | All | 3 | DCFace [22] |
ID R&D | USA | All | 4 | Novel |
K-IBS-DS | S. Korea | All | 5 | DCFace [22] | - |
CTAI | China | All | - |
Idiap-SynthDistill | Switzerland |
6 | Novel |
INESC-IGD |
All | 7 |
UNICA-IGD-LSI |
All | 8 | DCFace[22] |
SRCN_AIVL | China | 1.1 | 9 |
CBSR-Samsung | China |
10 | DCFace[22] |
BOVIFOCR-UFPR | Brazil |
11 | DCFace[22] | - |
@@
This team comprises members from Fudan University and Tencent Youtu Lab, China. They used a Latent Diffusion Model (LDM) [58] based on IDiff-Face [23] to synthesize faces. While IDiff-Face uses a noise embedding sampled from a Gaussian distribution to serve as the LDM’s, the ADMIS team uses identity embeddings as contexts, extracted from faces by a pretrained ElasticFace [71] model with an IResNet-101 [57] backbone. They trained the LDM with the CASIA-WebFace [51] database. As the LDM takes the ID embeddings as context, they considered an unconditional Denoising Diffusion Probabilistic Model (DDPM) trained on the FFHQ database [69] as a context generator. Specifically, they used the DDPM to generate faces with arbitrary identities, known as context faces. They exploited the same ElasticFace model to extract the embeddings from the context faces. To encourage the quality and the distinctiveness of identity from later LDM-generated faces, they filtered the embeddings by setting a minimum cosine similarity threshold of between arbitrary pairs of embeddings. This yields embeddings with discriminative identities. Furthermore, they accelerated the sampling process of the LDM by Denoising Diffusion Implicit Models (DDIM) [59]. For the training of the FR model, they generated images for each context. They adopted the ID oversampling strategy from DCFace [22] and performed it five times for each ID to enhance consistency. As a result, contexts were used for Sub-Tasks 1.1 and 2.1, while for Sub-Tasks 1.2 and 2.2. For Sub-Tasks 1.3 and 2.3, they expanded Sub-Tasks 1.1 and 2.1 with the CASIA-WebFace database. Both the baseline and proposed FR models used IResNet-101 architectures. They applied the ArcFace [6] loss with a batch size of and an initial learning rate of for epochs. The learning rate was divided by at epochs , , and . They also used random cropping augmentation during training. Their proposed architecture is described in Figure 2.
@@
This team comprises members from the Interactive Entertainment Group of Netease Inc., China. They initially used the DCFace [22] database, generating then more face images for each ID with large pose variations and occlusions using Photomaker [60]. Given a few input ID images, PhotoMaker can generate diverse personalized ID photos based on a text prompt while preserving the identity information from the input image. They randomly replaced these images in the original DCFace data to ensure that the total number of samples meets the requirement of . During the Photomaker inference, they adopted a batch size of and used random prompts including age, pose, and image quality to ensure the diversity of the generated samples. For Sub-Tasks 1.2 and 2.2, they combined this data with the version of DCFace, while for Sub-Tasks 1.3 and 2.3, it was merged with CASIA-WebFace [51]. For Sub-Tasks 1.2, 1.3, 2.2, and 2.3 they did not merge nor denoise samples from different databases, following the Partial FC approach [73], which consists in a sparse variant of model parallel architecture for training FR models. Regarding the FR model, they obtained the loss of different databases in independent AdaFace [7] heads, calculating the final loss as the average of the multiple heads. Both baseline and proposed models are based on IResNet-100 [57] architectures, with horizontal flipping. Their proposed architecture is described in Figure 3.
ID R&D (All sub-tasks)
@@
This team comprises members from ID R&D Inc, USA. To generate the synthetic data, they used two models trained on WebFace42M [74]. The first model was based on Hourglass Diffusion Transformers (HDT) [61], which combines the scalability of Transformer architectures with the efficiency of convolutional U-Nets [75]. It was trained focusing on conditional flow matching [76] and following the classifier-free guidance approach [77]. Identity embeddings were used directly, whereas style embeddings were processed through a Vector Quantised-Variational AutoEncoder (VQVAE) [62]. Specifically, for head position, embeddings were allocated for VQVAE processing, while age and facial expression attributes were represented with embeddings each. During inference, combinations of these embeddings were randomly selected. The second model used to generate synthetic data was based on StyleNAT [63], enhanced with a FR model. This network was trained using auxiliary sources of supervision: a pre-trained FR network with Prototype Memory [78] (for identity supervision) and a pre-trained face attribute classification network (for style supervision). To create their synthetic data, they used classifier weights of the trained Prototype Memory to get identity embeddings, of which were randomly selected and were uniformly sampled from the clusters obtained using k-means, to get demographic diversity. For each identity, they generated face images using each of the two generative models. In the first stage, identity embeddings were used by the HDT model to get images for each identity. Then of these images were included in the training dataset. Identity and style embeddings were taken from the remaining images and used as a condition to generate different images with the StyleNAT model. These images were also put into the training dataset. Regarding the FR model, they trained an IResNet-200 [57] with UniFace [65] loss for epochs. One network was trained with color, geometric augmentations, and FaceMix-B [64], and the other one using only random horizontal flipping. These two networks were combined in an “ensemble”, where the first one received the original image, and the second one a mirrored copy. They used the same model for Sub-Tasks 1.1, 1.2, 2.1 and 2.2. For Sub-Tasks 1.3 and 2.3, they combined the synthetic data with the CASIA-WebFace [51] data, training two models, one on the mixed data, and the other on the CASIA-WebFace. Their proposed architecture is described in Figure 4.
@@
This team comprises members from the Korea Advanced Institute of Science & Technology and the Institute for Basic Science, South Korea. They used DCFace [22] with and face images (depending on the sub-task). Regarding the FR model, they used several IResNet [57] models of , , and layers with Squeeze-and-Excitation blocks [67], which are architectural components designed to enhance the representational power of convolutional neural networks by dynamically adjusting channel-wise features. Inspired by SlackedFace [66], they made two modifications to enhance the AdaFace [7] FR classifier: i) they used renormalized uniform initialization as a more reliable weight initialization for uniformity across identity prototypes in the unit sphere and ii) replaced the L2-norm with the powered-norm (p-norm), or face recognizability index from [66], which integrates the L2-norm with the learned embedding proximity. The training stage was in line with [22] and [29], including optimizer, learning rate, etc. For Sub-Tasks 1.3 and 2.3, the first subjects of the CASIA-WebFace [51] were assigned for training, and the remaining ones for performance validation using random pairs with challenging conditions (identified based on the poorest L2-norm values [21]). The final score was obtained by aggregating the comparison scores of the different IResNet models, along with the horizontally flipped instances through score fusion. All training and test sets were realigned using RetinaFace [79], followed by a similarity transformation. For all sub-tasks, aggressive data augmentations were applied, including random horizontal flipping, photometric operations, cropping, resizing, and the addition of sunglasses and masks. Their proposed architecture is described in Figure 5.
@@
This team comprises members from China Telecom AI, China. By analyzing popular synthetic data, they found that intra-class and inter-class noise was widely present. Data cleaning can effectively remove the bad examples of synthetic data and retain important images from a large amount of synthetic data. In order to select the optimal synthetic data, they first trained an IResNet-100 [57] model with Squeeze-and-Excitation blocks [67] using CASIA-WebFace [51] to extract features of synthetic images from DCFace [22], GANDiffFace [21], and DigiFace-1M [29]. Subsequently, they used the DBSCAN clustering method to segregate intra-class noise and removed IDs with a class center feature cosine similarity greater than . Finally, they used the cleaned synthetic data merged with CASIA-WebFace to finetune the IResNet-100 for a second data refinement. From the final refined synthetic dataset, they sampled face images while retaining as many IDs as possible to build their synthetic training set. Regarding the FR model, in particular Sub-Task 2.3 in which they achieved their highest position among all sub-tasks, they trained IResNet-100 with AdaFace loss [7] (A1) and CosFace loss [68] (A2) with mask and occlusion augmentation on CASIA-WebFace and the refined synthetic data. They used an ensemble of A1, A2, and a model trained with only synthetic data. Furthermore, data augmentation was considered to enhance all features.
Idiap-SynthDistill (Sub-Tasks 1.2 and 2.2)
@@
This team comprises members from the Idiap Research Institute, EPFL, and Université de Lausanne, Switzerland. The proposed method was based on SynthDistill [70], which is an end-to-end approach, generating synthetic images and training the FR model in the same training loop. Instead of using the pre-trained model in a separate step, they directly used it in the training loop for supervision, while a new student FR model was trained fully using synthetic data generated from a StyleGAN model [69]. For generating synthetic images, they trained StyleGAN2 [80] with the CASIA-WebFace database [51] and then dynamically generated 20 synthetic images during training based on the training loss. For the dynamic image generation, they used the training loss from every iteration as feedback to find the most difficult synthetic image in each batch and then re-sampled a new batch of synthetic images in the intermediate latent space of StyleGAN near the latent vector of the most difficult sample. If the loss value was high, they re-sampled with a relatively small standard deviation around the difficult sample and generated similar images, but if the loss value was small they re-sampled with a higher standard deviation, generating images with more variations. Throughout the process, the generated images were resized to before being fed to the FR models. They used a pre-trained FR model with IResNet-101 [57] architecture trained with CosFace [68] loss on a subset of the WebFace260M database [74] and trained a new model as a student network with the same architecture using synthetic data with their dynamic synthetic image generation approach. They used the Adam optimizer with an initial learning rate of and trained their student model with the same loss function as in [70]. For thresholding, a subset of DCFace [22] was used to determine the optimal threshold for maximizing verification accuracy, using a 10-fold cross-validation approach based on a random selection of identities and comparison pairs. Their proposed architecture is described in Figure 6.
@@
UNICA-IGD-LSI (All sub-tasks)
This team comprises members from Fraunhofer IGD, Germany, University of Cagliari, Italy, and University of Ljubljana, Slovenia. They used the DCFace [22] synthetic database as it led to remarkable performance gains under well-known evaluation benchmarks for face verification, while combined with real data [84]. Also, they considered synthetic data generated with IDiff-Face [23] and ExFaceGAN [72] in Sub-Tasks 1.1, 1.2, 2.1, and 2.2. The ExFaceGAN data was generated using an identity disentanglement approach on pretrained GAN-Control [85]. Regarding the FR model, they trained a ResNet-100 [34] network using CosFace loss [68] with a margin penalty of 0.35 and a scale term of 64. The similarity mean difference between real-only and synthetic-only samples () was scaled and added to the loss value. They trained the FR model for epochs with a batch size of and an initial learning rate of , which was divided by after , , , and epochs. During the training phase, the synthetic samples were augmented using RandAugment with operations and a magnitude of 16, following [86, 84]. For Sub-Tasks 1.3 and 2.3, the selected synthetic dataset was combined with CASIA-Webface [51], obtaining a total of images from identities. Their proposed architecture is described in Figure 8.